Four kg of water is placed in an enclosed volume of 1m^3. Heat is added until the temperature is 150∘C. Find

( a) the pressure,

(b )the mass of vapor, and

(c) the volume of the vapor.


1
Expert's answer
2021-11-05T03:40:00-0400

(a)

From steam tables,

P=476.16kPa

(b)

"v=v_f+x(v_g-v_f)"

"0.25=0.001091+x(0.39248-0.001091)"

0.25=0.001091+0.391389x

0.248990=0.391389x

"x=\frac{0.248990}{0.391389}"

x=0.63596

"x=\frac{m_g}{m}"

"m_g=x*m"

=0.63596*4

=2.544kg

(c)

"V_g=v_g*m_g"

=0.39248*2.544

=0.998m3